OpenStudy (anonymous):

Write the fraction as a percent 15/9

OpenStudy (nottim):

Errr...What grade are you?

hero (hero):

\[\frac{is}{of} = \frac{\%}{100} \]

OpenStudy (nottim):

yup yup.

hero (hero):

\[\frac{15}{9} = \frac{x}{100}\]

OpenStudy (nottim):

Yup yup yup.

hero (hero):

9x = 1500 x = 1500/9 x = 166.67 percent

OpenStudy (nottim):

Yup yup yup yup.